OpenStudy (anonymous):

Which of the following is equivalent to log (base 6) 37 rounded to three decimal places? a. 1.677 b. 0.496 c. 0.790 d. 2.015 I would use a graphing calculator but I loaned it to my friend.

OpenStudy (campbell_st):

well you can use change of base formula to evaluate it \[\log_{a}(b) = \frac{\ln(b)}{\ln(a)}\] hope it helps

OpenStudy (campbell_st):

but here is a hint.... whats the value of 6^2

OpenStudy (anonymous):

36

OpenStudy (campbell_st):

so remember.... the value of a log is the power of the base.... so using the hint... the value is a bit more than 2....

OpenStudy (anonymous):

2.015?

OpenStudy (campbell_st):

to check... whats the value of 6^2.015)

OpenStudy (anonymous):

I got 36.98

OpenStudy (campbell_st):

thats it... a little intuition can eliminate the need for calculations

OpenStudy (anonymous):

:O I understand! Haha thank you so much! :D